OpenStudy (anonymous):

List the critical numbers of the following function in increasing order: g(x)=3x^(1/3)-3x^(-2/3)

OpenStudy (anonymous):

\[g(x)= 3x ^{1/3}-3x ^{-2/3}\]

OpenStudy (bahrom7893):

Have you ever done derivatives?

OpenStudy (anonymous):

yeah i got the derivative. And then i know i have to set it to 0. but then i'm having problems solving for x

OpenStudy (bahrom7893):

Ok, what's your derivative?

OpenStudy (anonymous):

\[g'(x)=x ^{-2/3}+2x ^{-5/3}\]

OpenStudy (anonymous):

Then when I set it equal to 0 and try to solve, the fraction exponents are messing with me hah

OpenStudy (bahrom7893):

Rewrite this in fractional form:\[g'(x)=\frac{1}{x^{2/3}}+\frac{2}{x^{5/3}}\]Find the common denominator. What do you get.

OpenStudy (anonymous):

\[\frac{ x ^{5/3}+2x ^{2/3} }{ x ^{10/9} }\]

OpenStudy (anonymous):

sorry denominator should be x^(-7/3)

OpenStudy (anonymous):

x^(7/3)! my bad haha

OpenStudy (bahrom7893):

Not exactly, it's just algebra at this point, so I'll help you out. This is what that results in when combined:\[\frac{x+2}{x^{5/3}}\]

OpenStudy (bahrom7893):

Actually you just fixed your error... but when simplified it would amount to my answer anyway.

OpenStudy (anonymous):

Oh yeah, cool! i can solve it from there then. Thanks man!

OpenStudy (bahrom7893):

Btw, don't forget. You have a critical point when the derivative is undefined as well. So don't forget to solve:\[x^{5/3}=0\]as well.

OpenStudy (anonymous):

That would just be 0 wouldn't it?

OpenStudy (anonymous):

so -2 and 0 would be the c.p.'s?

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

Thanks

OpenStudy (bahrom7893):

np